OpenStudy (anonymous):

how much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 60% antifreeze?

OpenStudy (jamesj):

Let X be the total volume of the final mixture. Then you know that 60% of X is antifreeze. Hence \[ 0.6X = 1 \ gallon \] or \[ X = \frac{1}{0.6} \ gallons \] Now can you find the amount of water you need to add?

OpenStudy (jamesj):

Notice that \( X = 1 \frac{2}{3} \) gallons. Hence how much water is added?

OpenStudy (anonymous):

i'm sorry i'm totally clueless....i feel like x should be the amount of water?

OpenStudy (jamesj):

X is the volume of the final mixture.

OpenStudy (jamesj):

And you want 60% of X to be the anti-freeze. Now 60% of X is \[ \frac{60}{100} X \] So far, so good?

OpenStudy (anonymous):

so then x-1/.6 is the water?

OpenStudy (jamesj):

No.

OpenStudy (anonymous):

40 percent...?

OpenStudy (jamesj):

X = (antifreeze) + (water)

OpenStudy (jamesj):

And we want that 60% 40% X = (anti-freeze) + (water)

OpenStudy (jamesj):

But we also are given 1 gallon of anti-freeze. Therefore it must be that 60% of X = 1 gallon. Yes?

OpenStudy (anonymous):

right

OpenStudy (jamesj):

Now 60% of X is \[ \frac{60}{100}X \] if that is equal to 1, then \[ X = \frac{100}{60} = \frac{5}{3} \] ok?

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

Now because X = (anti-freeze) + (water) and X = 5/3 what must the volume of water be?

OpenStudy (anonymous):

1-5/3?

OpenStudy (jamesj):

X = (anti-freeze) + (water) i.e., 5/3 = 1 + (water) i.e., (water) = 5/3 - 1 gallons = 2/3 gallons

OpenStudy (anonymous):

ohhh ok, man i am so bad at word problems

OpenStudy (jamesj):

There's another way to solve this using proportions.

OpenStudy (jamesj):

If 60% of the solution is antifreeze, then 40% is water and it must be that \[ \frac{antifreeze}{water} = \frac{60}{40} \] Now, as the volume of anti-freeze is 1 gallon, we have that \[ \frac{1}{water} = \frac{60}{40} \] Now solve for Water.

OpenStudy (anonymous):

oh i think that one makes more sense to me

OpenStudy (anonymous):

can i ask you another one or should i post it on the board again?

OpenStudy (jamesj):

post a new question.